# Just Trial Once: Ongoing Causal Validation of Machine Learning Models

### v1

@@ Appendix B Proofs
11 11 $=\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{X\in\mathcal{X}_{\not\geq},\pi_{e}(X)\neq a_{0}\}$
$\quad+\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}+\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}$ (36)
and the argument follows similarly, such that the upper bound follows from the application of Sections B.1 and B.1 to each term.
First, $\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{X\in\mathcal{X}_{=},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{=}(x)]$ follows from Section B.1.
Next, $\mathbf{1}\{X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{>}(x)]$ follows from Section B.1. Alternatively, $\mathbf{1}\{X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{X\in\mathcal{X}_{>},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,D\in\tilde{\mathcal{S}}_{>}(x)]$, where we replace $\mathcal{S}_{>}(x)$ with $\tilde{\mathcal{S}}_{>}(x)$, is also true from Section B.1.
Next, $\mathbf{1}\{X\in\mathcal{X}_{\not\geq},\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{X\in\mathcal{X}_{\not\geq},\pi_{e}(X)\neq a_{0}\}Y_{max}$ follows from Section B.1.
Next, $\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{X\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,D\in\mathcal{S}_{*}(x)]$ follows from Section B.1.
Finally, $\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{X\not\in\mathcal{X}_{*},\pi_{e}(X)=a_{0}\}Y_{max}$ follows from Section B.1. ∎
@@
We will start by observing that certain terms cancel out in the difference of the bound $U(\pi_{e})-L(\pi_{e})$. We begin by recalling the definition of these bounds from Assumption 3.3
$L(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]$ (37)
$\quad+\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$ (38)
$\quad+\mathbf{1}\{x\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{min}]$ (39)
$U(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{=}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{=}(x)]$ (40)
$\quad+\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$
$\quad+\mathbf{1}\{x\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{*}(x)]$ (41)
$\quad+\mathbf{1}\{x\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{max}]$ (42)
By linearity of expectation, we can remove identical terms, i.e., we can observe that in the difference $U(\pi_{e})-L(\pi_{e})$, the terms in Eqs. 37 and 40 cancel, and likewise the terms in Eqs. 38 and 41 cancel, leaving us with the following after collecting similar terms Eqs. 39 and 42
@@
$=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$ (43)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$ (44)
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ (45)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ (46)
$\qquad+(\mathbf{1}\{x\not\in\mathcal{X}_{*}\}\mathbf{1}\{\pi_{e}=a_{0}\})(Y_{\text{max}}-Y_{\text{min}})]$ (47)
@@
  • First, we note that we can partition $\mathcal{X}_{<}$ into two subsets: The set of $x$ where the only agreeing policies are those with worse performance, and the set where there also exist policies with greater performance (note that the existence of policies where performance is equivalent is ruled out by the definition of $\mathcal{X}_{<}$). We can argue similarly for $\mathcal{X}_{>}$, and write
    $\mathbf{1}\{x\in\mathcal{X}_{>}\}$ $=\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}+\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}\}$
    $\mathbf{1}\{x\in\mathcal{X}_{<}\}$ $=\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}+\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}\}$
  • Second, we note that we can partition $\mathcal{X}_{\not\geq}=\mathcal{X}_{*}^{c}\cup\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}$, and likewise $\mathcal{X}_{\not\leq}=\mathcal{X}_{*}^{c}\cup\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}$. For instance, the set of $x$ where there are no trialed policies (with equal or better performance than $\pi_{e}(x)$) that agree with $\pi_{e}(x)$ is equal to the set of $x$ where there are no agreeing policies at all ($\mathcal{X}_{*}$) plus the set of $x$ where the only agreeing policies have worse performance ($\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}$). As a result, we can write that
    $\mathbf{1}\{x\in\mathcal{X}_{\not\leq}\}$ $=\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}+\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}$
    $\mathbf{1}\{x\in\mathcal{X}_{\not\geq}\}$ $=\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}+\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}$
@@
$=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$ From Eq. 43 (48)
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]$ From Eq. 43 (49)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$ From Eq. 44 (50)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)]$ From Eq. 44 (51)
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ From Eq. 45 (52)
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ From Eq. 45 (53)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ From Eq. 46 (54)
$\qquad-\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ From Eq. 46 (55)
$\qquad+(\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}\mathbf{1}\{\pi_{e}=a_{0}\})(Y_{\text{max}}-Y_{\text{min}})]$ From Eq. 47 (56)
@@
$=\mathbb{E}[\mathbf{1}\{x\in\mathcal{X}_{*}^{c}\}(Y_{\text{max}}-Y_{\text{min}})$ Eqs. 53, 55 and 56
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}(\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]-\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)])$ Eqs. 48 and 50
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{<}\cap\mathcal{X}_{>}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}(Y_{\text{max}}-\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)])$ Eqs. 51 and 52
$\qquad+\mathbf{1}\{x\in\mathcal{X}_{>}\cap\mathcal{X}_{<}^{c}\}\mathbf{1}\{\pi_{e}\neq a_{0}\}(\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]-Y_{\text{min}})$ Eqs. 49 and 54
Which gives us the desired result of the form $\mathbb{E}[\delta(X,Y,D)]$. The fact that $\delta(X,Y,D)\geq 0$ almost surely follows from the fact that $Y_{\text{max}}\geq\mathbb{E}[Y\mid C]\geq Y_{\text{min}}$ for any conditioning set $C$, and the fact that $(\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{>}(x)]-\mathbb{E}[Y\mid X=x,D\in\mathcal{S}_{<}(x)])$ is nonnegative by Assumption 3.1. Because $L(\pi_{e})$ and $U(\pi_{e})$ may be alternatively expressed with $\tilde{\mathcal{S}}_{<}(x)$ and $\tilde{\mathcal{S}}_{>}(x)$, $\delta(X,Y,D)$ may be expressed with $\tilde{\mathcal{S}}_{<}(x)$ and $\tilde{\mathcal{S}}_{>}(x)$ as well. ∎
@@
$\tilde{\mathcal{S}}_{<}(x)$ $=\operatorname*{arg\,max}_{i\in\mathcal{S}_{<}(x)}f_{M}(\pi_{i})$ $\tilde{\mathcal{S}}_{>}(x)$ $=\operatorname*{arg\,min}_{i\in\mathcal{S}_{>}(x)}f_{M}(\pi_{i}).$
@@
Shared Structural Equations for $\mathcal{M}_{L},\mathcal{M}_{U}$: Both SCMs share the following equations that respect Assumptions 3.1 and 3.2, which give rise to a shared distribution over $P(D,X,M,A)$, and these can be chosen to match any such observed distribution.
@@
    $\mathcal{S}_{a}(x)\coloneqq\{i\in\mathcal{D}\mid\pi_{i}(x)=a\}$
    be the set of all policies (possibly an empty set) that have output $a$ on the input $x$, and let $\mathbf{M}_{a}(x)=(m_{1},\ldots,m_{K})$ be the (ordered) set of performance values for these policies, where $m_{1}$ denotes the worst performance $f_{M}(\pi_{i})$ (breaking ties arbitrarily) of all policies $\pi_{i}$ where $i\in\mathcal{S}_{a}(x)$, and $m_{K}$ denotes the best performance, where $K$ is the size of the set $\mathcal{S}_{a}(x)$.
@@
    • If $\mathcal{S}_{a}(x)$ is empty for this $a$, then for all $m\in\mathbb{R}$, we let $f_{Y}^{L}(x,a,m,\epsilon_{Y})=Y_{\text{min}}$, $f_{Y}^{U}(x,a,m,\epsilon_{Y})=Y_{\text{max}}$. In other words, at this point in $x$, if no trialed policy takes action $a$, we assume the worst (for $f_{Y}^{L}$) and the best (for $f_{Y}^{U}$) possible outcomes. We can easily verify that both $f_{Y}^{L},f_{Y}^{U}$ satisfy Assumptions 3.1 and 3.2 since the output is a constant for each function, and satisfy Assumption 3.3 since the output remains bounded.
    • If $\mathcal{S}_{a}(x)$ (and consequently $\mathbf{M}_{a}(x)$) is non-empty, then we first define the behavior of $f_{Y}^{L},f_{Y}^{U}$ at all the (observable) performance values in $\mathbf{M}_{a}(x)$ to match the conditional distribution $P(Y\mid x,a,m)$.

### v2

@@ Appendix D Proofs
11 11 $=\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}$ (37)
and the argument follows similarly, such that the upper bound follows from the application of Sections D.1 and D.1 to each term.
First, $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)]$ follows from Section D.1.
Next, $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\pi_{e}(X)\neq a_{0}\}Y_{max}$ follows from Section D.1.
Next, $\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]=\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,\Pi\in\mathbf{\Pi}^{e}(x)]$ follows from Section D.1.
Finally, $\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}\mathbb{E}[Y\mid X,A=\pi_{e},M=f_{M}(\pi_{e})]\leq\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing,\pi_{e}(X)=a_{0}\}Y_{max}$ follows from Section D.1. ∎
@@
We will start by observing that certain terms cancel out in the difference of the bound $U(\pi_{e})-L(\pi_{e})$. We begin by recalling the definition of these bounds from Remark 3.1, rearranged slightly.
$L(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$
$\quad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$ (38)
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{min}]$ (39)
$U(\pi_{e})$ $=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$
$\quad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\mathbf{\Pi}^{e}(x)]$ (40)
$\quad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\}Y_{max}]$ (41)
By linearity of expectation, we can remove identical terms, i.e., we can observe that in the difference $U(\pi_{e})-L(\pi_{e})$, the terms in Eqs. 38 and 40 cancel, leaving us with the following after collecting similar terms Eqs. 39 and 41.
@@
$=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$ (42)
$\qquad-\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$ (43)
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ (44)
$\qquad-\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ (45)
$\qquad+(\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\})(Y_{\text{max}}-Y_{\text{min}})]$ (46)
@@
  • First, we note that we can partition the subset of $\mathcal{X}$ satisfying $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ into two further subsets: The set of $x$ where the only agreeing policies are those with worse performance, and the set where there also exists agreeing policies with equal or greater performance. Note that, if there exists trial policies with exactly equal performance to the new policy, both $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$ and $\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}$ must be true. We can argue similarly for $\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}$, and write
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}=\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}$
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}=\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$
  • Second, we note that we can partition the subset of $\mathcal{X}$ satisfying $\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}$ into two further subsets: The set of $x$ where the only agreeing trial policies have greater performance, and the set where there are no agreeing trial policies at all. We can argue similarly for $\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}$, and write
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}=\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}$
    $\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}=\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}$
@@
$=\mathbb{E}[\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$ From Eq. 42 (47)
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]$ From Eq. 42 (48)
$\qquad-\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$ From Eq. 43 (49)
$\qquad-\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)]$ From Eq. 43 (50)
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ From Eq. 44 (51)
$\qquad+\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{max}$ From Eq. 44 (52)
$\qquad-\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ From Eq. 45 (53)
$\qquad-\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}Y_{min}$ From Eq. 45 (54)
$\qquad+(\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}\mathbf{1}\{\pi_{e}=a_{0}\})(Y_{\text{max}}-Y_{\text{min}})]$ From Eq. 46 (55)
@@
$=\mathbb{E}[\mathbf{1}\{\mathbf{\Pi}^{e}(X)=\varnothing\}(Y_{\text{max}}-Y_{\text{min}})$ Eqs. 52, 54 and 55
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}$
$\qquad\qquad(\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]-\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)])$ Eqs. 47 and 49
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)=\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)\neq\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}(Y_{\text{max}}-\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)])$ Eqs. 50 and 51
$\qquad+\mathbf{1}\{\tilde{\mathbf{\Pi}}^{e}_{\geq}(X)\neq\varnothing,\tilde{\mathbf{\Pi}}^{e}_{\leq}(X)=\varnothing\}\mathbf{1}\{\pi_{e}\neq a_{0}\}(\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]-Y_{\text{min}})]$ Eqs. 48 and 53
After rearranging terms slightly, this gives us the desired result of the form of $\delta(X,Y,\Pi)$ stated in the theorem. The fact that $\delta(X,Y,\Pi)\geq 0$ almost surely follows from the fact that $Y_{\text{max}}\geq\mathbb{E}[Y\mid C]\geq Y_{\text{min}}$ for any conditioning set $C$ according to Assumption 3.3, and the fact that $(\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)]-\mathbb{E}[Y\mid X=x,\Pi\in\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)])$ is nonnegative by Assumption 3.1. ∎
@@
$\tilde{\mathbf{\Pi}}^{e}_{\leq}(x)$ $=\operatorname*{arg\,max}_{\pi\in\mathbf{\Pi}^{e}_{\leq}(x)}f_{M}(\pi)$ $\tilde{\mathbf{\Pi}}^{e}_{\geq}(x)$ $=\operatorname*{arg\,min}_{\pi\in\mathbf{\Pi}^{e}_{\geq}(x)}f_{M}(\pi).$
@@
Shared Structural Equations for $\mathcal{M}_{L},\mathcal{M}_{U}$: Both SCMs share the following equations that respect Assumptions 3.1 and 3.2, which give rise to a shared distribution over $P(X,A,M,\Pi,D)$, and these can be chosen to match any such observed distribution.
@@
    ${\bf\Pi}^{a}(x)\coloneqq\{\pi\in\Pi\mid\pi(x)=a\}$
    be the set of all policies (possibly an empty set) that have output $a$ on the input $x$, and let $\mathbf{M}_{a}(x)=(m_{1},\ldots,m_{K})$ be the (ordered) set of performance values for these policies, where $m_{1}$ denotes the worst performance $f_{M}(\pi_{i})$ (breaking ties arbitrarily) of all policies $\pi_{i}$ where $\pi_{i}\in{\bf\Pi}^{a}(x)$, and $m_{K}$ denotes the best performance, where $K$ is the size of the set ${\bf\Pi}^{a}(x)$.
@@
    • If ${\bf\Pi}^{a}(x)$ is empty for this $a$, then for all $m\in\mathbb{R}$, we let $f_{Y}^{L}(x,a,m,\epsilon_{Y})=Y_{\text{min}}$, $f_{Y}^{U}(x,a,m,\epsilon_{Y})=Y_{\text{max}}$. In other words, at this point in $x$, if no trialed policy takes action $a$, we assume the worst (for $f_{Y}^{L}$) and the best (for $f_{Y}^{U}$) possible outcomes. We can easily verify that both $f_{Y}^{L},f_{Y}^{U}$ satisfy Assumptions 3.1 and 3.2 since the output is a constant for each function, and satisfy Assumption 3.3 since the output remains bounded.
    • If ${\bf\Pi}^{a}(x)$ (and consequently $\mathbf{M}_{a}(x)$) is non-empty, then we first define the behavior of $f_{Y}^{L},f_{Y}^{U}$ at all the (observable) performance values in $\mathbf{M}_{a}(x)$ to match the conditional distribution $P(Y\mid x,a,m)$.